The foci of the hyerbola are S(-3,-2) and S`(5,6). If it`s e=2 then the equation of it`s directrix corresponding to focus S is

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
a=-3/2
B^2 = a^2(1-e^2) = 3a^2 = 27/4
     B= (3√3)/2
Point on directrix = (-2,3√3/2 ) ; slope of directrix = -1 (perpendicular to axis)
Equation : x+y+2-3√3/2 =0 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)